# SPADE: Sparsity-Guided Debugging for Deep Neural Networks

### v1

@@ B.2 MobileNet
Table B.3: MobileNet model results. Sparsity ratios tuned using ImageNet model. ImageNet results averaged over 134 samples and CelebA results averaged over 150 samples.
Saliency Method ImageNet CelebA AUC Pointing Game AUC Pointing Game Dense SPADE $\Delta$ Dense SPADE $\Delta$ Dense SPADE $\Delta$ Dense SPADE $\Delta$ Saliency 88.9 93.04 +4.14 93.23 94.03 +0.8 95.43 96.92 +1.49 80.67 80.0 -0.67 DeepLift 85.71 90.7 +4.99 81.34 81.34 0.0 93.26 96.15 +2.89 70.0 81.33 +11.33 Guided Backprop 88.91 93.04 +4.12 93.28 94.03 +0.75 95.43 96.92 +1.49 80.67 83.33 +2.66 Guided Grad-Cam 95.19 95.73 +0.54 93.28 94.78 +1.5 86.76 86.85 +0.1 66.0 68.0 +2.0 Lime 89.45 91.62 +2.16 70.15 68.66 -1.49 67.64 77.14 +9.5 51.33 64.67 +13.34 Occlusion 89.51 90.98 +1.47 94.03 94.03 0.0 90.39 94.66 +4.28 83.33 95.33 +12.0 Integrated Gradients 89.76 92.88 +3.12 87.22 88.06 +0.84 95.91 97.79 +1.88 76.67 79.33 +2.66 Gradient Shap 89.45 92.07 +2.62 84.96 84.33 -0.63 93.94 96.24 +2.3 76.67 76.67 0.0 Average 89.61 92.51 +2.90 87.19 87.41 +0.22 89.84 92.83 +2.99 73.17 78.58 +5.41
B.3 ConvNext
We additionally conducted ImageNet and CelebA experiments on the ConvNext-T (Liu et al., 2022) architecture. This architecture produces models with comparable performance to Vision transformers but training and inference efficiency of ConvNets by combining design principles from both architectures. Similar to MobileNet, we exclude depthwise covolutions and only prune pointwise convolutions and linear layers. As with MobileNet, we omit LRP from this analysis, due to unspecified behaviour for this method in cases where non-ReLU (here, GeLU activations) are used, and, like with MobileNet, we combine the InputXGradient and DeepLift rows. For this architecture, Gaussian Noise and Random Masking were added to the image augmentations. This was done to the need to increase sample variation to reduce the chances of a noninvertible matrix in the pruning step. The augmented samples may be seen in Figure C.2.
@@ B.3 ConvNext
The results are presented in Table B.4. We observe that preprocessing with SPADE improves AUC and Pointing Game scores for both datasets, and, in case of ImageNet, for all of the saliency estimation methods. On average, SPADE preprocessing improves ImageNet Saliency AUC by 3.09% and pointing game accuracy by 5.1%. On CelebA, SPADE improves ImageNet saliency AUC by 1.38% and Pointing Game AUC by 1.87%.
@@
Table B.4: ConvNext-T model results. Sparsity ratios tuned using ImageNet model. ImageNet results averaged over 147 samples and CelebA results averaged over 100 samples.
@@
Saliency Method ImageNet CelebA AUC Pointing Game AUC Pointing Game Dense SPADE $\Delta$ Dense SPADE $\Delta$ Dense SPADE $\Delta$ Dense SPADE $\Delta$ Saliency 85.24 87.5 +2.25 82.31 85.03 +2.72 96.6 96.95 +0.35 76.0 75.0 -1.0 DeepLift 81.95 84.6 +2.64 71.43 80.27 +8.84 94.93 95.53 +0.6 59.0 64.0 +5.0 Guided Backprop 85.24 87.5 +2.25 84.35 85.03 +0.68 96.6 96.95 +0.35 76.0 75.0 -1.0 Guided Grad-Cam 84.1 91.99 +7.89 82.99 88.44 +5.45 87.05 90.19 +3.13 73.0 79.0 +6.0 Lime 93.41 94.73 +1.32 70.75 75.51 +4.76 75.3 73.78 -1.53 59.0 58.0 -1.0 Occlusion 85.27 88.24 +2.97 87.76 88.44 +0.68 89.53 92.2 +2.67 83.0 88.0 +5.0 Integrated Gradients 87.0 89.13 +2.12 71.43 82.31 +10.88 92.76 95.55 +2.79 64.0 67.0 +3.0 Gradient Shap 86.2 89.44 +3.24 72.79 79.59 +6.8 91.71 94.36 +2.66 64.0 63.0 -1.0 Average 86.05 89.14 +3.09 77.98 83.08 +5.1 90.56 91.94 +1.38 69.25 71.12 +1.87
Appendix C Additional hyperparameters
@@
Table C.5: ImageNet Trojan patches with their source and target class. ”Any” means any image could be used for the Trojan. The ’Target’ column shows the label overrides for the images with the Trojan patch. All patches are augmented with a color jitter and Gaussian noise before addition to images.
@@
Table C.6: CelebA Trojan patches. All images may be chosen for a Trojan. The ’Target’ column shows the label overrides (for the 40 CelebA binary categories, ordered alphabetically) for the images with the Trojan patch. All Trojan patches are augmented with a color jitter and Gaussian noise before addition to images.
@@
Table C.7: Food-101 Trojan patches with their source and target class. ”Any” means any image could be used for the Trojan. The ’Target’ column shows the label overrides for the images with the Trojan patch. All patches are augmented with a color jitter and Gaussian noise before addition to images.
@@
Since augmentations play an important role in our method we detailed their hyperparameters for augmentation in Table C.8. We also show typical augmented samples in Figure C.1, and Figure C.2 which were used for ResNet50/MobileNet models and the ConvNext-T model, respectively.
Table C.8: Augmentation details. “Models” column explain which models used the augmentation. Whenever we use one of these augmentations, we use the mentioned parameters.
@@
Guassian Noise $\sigma^{2}$ = 0.001 ConvNext
@@
When training ResNet50 on Food-101 dataset we used the hyperparameters suggested in Kornblith et al. (2019), with includes a weight decay of 0.0005. Other hyperparameters are highlighted in Table C.9.
For other cases which includes ResNet50, MobileNet, or ConvNext-T on ImageNet, or celebA dataset, we use a 0.9 momentum and step-lr learning rate scheduler with a step-lr-gama 0.1 for all backdoorings and a weight decay of 0.0001. The initial learning rate is chosen from the options - 0.01, 0.001, 0.0001, 0.00001 - based on accuracy on Trojan samples at the end of training. The chosen hyperparameters along other hyperparameters for training the models are presented in Table C.9.
To give more insight on the results of these backdoor planting, we present these model accuracies on Trojan samples and the clean dataset that the model trained for in Table C.10. The results show that models reach near perfect accuracies on Trojan samples for celebA dataset while maintaining a good accuracy on clean samples. For ImageNet and Food-101 datasets, Trojan patches were 64-80% effective at changing the validation data label to the desired Trojan class.
@@
Figure C.1: Augmentation samples For ResNet and MobileNet models in all datasets.
@@
Figure C.2: Augmentation samples For ConvNext model
Table C.9: Hyperparameters used for planting backdoors in the models.”Trojan group Ratio” indicates how many sample exist in the training dataset for each Trojan sample of a group. ”step-lr” refers to the epoch that learning rate drops.
Model DataSet Trojan group Ratio Batch Size Learning Rate step-lr Epoch
@@
Table C.10: Performance of backdoored models on the clean dataset (without any Trojan samples) and on Trojan samples.
@@
Appendix D Ablation study
In this section, we examine how the various hyperparameters of SPADE that impact its performance on the saliency map accuracy task.
D.1 Sample selection
We first investigate the impact of varying the sample size and selection for the Optimal Brain Damage (OBD) pruning process. We experimented with different sample selection methods, namely:
@@ D.1 Sample selection
Table D.11: Impact of sample selection for the network pruning step of SPADE. 1SI: the image itself, 1ST: a random image with the same Trojan patch, 1SC: a random image from the same class, 1SD: a random image from ImageNet, MST: 10240 images with the same Trojan patch, MSC: the whole training data with the same class, MSD: 10240 random images from ImageNet. Based on 100 samples. The First number in each cell refers to AUC and the Second number refers to Point Game measure.
Saliency Method Dense 1SI 1ST 1SC 1SD MST MSC MSD saliency 86.5/76 95.2/87 60.8/32 46.5/6 48.0/11 60.3/28 41.0/5 43.4/4 InputXGradient 82.8/60 92.9/82 60.0/22 50.2/5 50.1/4 59.0/18 50.0/6 50.2/6 DeepLift 93.0/81 94.7/82 60.3/21 50.9/6 50.2/7 57.5/10 50.7/7 50.8/3 LRP 92.1/66 99.1/99 83.6/46 77.6/25 81.3/36 84.3/49 72.9/21 72.8/25 Guided Backprop 95.3/94 96.9/93 83.1/57 76.4/35 80.8/55 83.8/59 70.9/20 77.2/42 Guided Grad-Cam 97.8/95 98.1/93 83.6/58 71.3/32 70.3/46 84.9/57 67.0/16 65.2/39 Lime 92.7/74 95.6/74 74.7/40 61.3/31 53.1/16 75.5/44 63.4/31 52.0/19 Occlusion 86.1/92 94.6/92 65.7/42 48.5/11 54.8/12 68.0/41 43.8/6 48.2/7 IntegratedGradients 87.5/69 94.5/80 62.4/22 50.3/5 51.9/9 60.3/17 50.2/3 50.2/5 gradientSHAP 87.2/69 94.4/80 62.4/22 50.2/6 52.1/11 60.3/18 50.1/4 50.2/4 Average 90.1/77.6 95.6/86.2 69.7/36.2 58.3/16.2 59.3/20.7 69.4/34.1 56.0/11.9 56.0/15.4
@@
The results, summarized in Table D.11, show clearly that the use of the single, augmented sample for the pruning step of SPADE is crucial for the efficacy of the method. More generally, using images with the same Trojan patch yielded better results than other sample selection methods, while using images with the same base class was no better than using randomly chosen images from the entire dataset. Further, this demonstrates that the act of pruning alone does not necessarily enhance interpretability. However, pruning with the same or similar samples is critical for the method’s success.
D.2 Choice of augmentation

### v2

@@ B.3 Transferability of Layer Sparsity Targets across Datasets
Table 5: ImageNet, ResNet transferability of sparsity ratio over datasets. The sparsity ratios were tuned using ImageNet and used in these experiments. The results averaged over 100 samples for each of these datasets and interpretability methods.
Saliency Method CelebA (ImageNet Pretrained) Food101 (Random Initialization) Dense SPADE $\Delta$ Dense SPADE $\Delta$ Saliency 73.52 92.81 +19.28 69.13 94.62 +25.49 InputXGradient 68.26 92.09 +23.84 66.09 93.48 +27.39 DeepLift 87.76 91.21 +3.45 89.41 95.18 +5.77 LRP 86.82 96.8 +9.98 87.26 98.64 +11.38 GuidedBackprop 97.87 96.63 -1.24 98.26 98.44 +0.18 GuidedGradCam 88.89 89.13 +0.24 97.57 97.61 +0.03 Lime 75.58 62.42 -13.16 91.76 93.66 +1.9 Occlusion 65.12 79.27 +14.15 75.87 91.45 +15.58 IntegratedGradients 83.01 93.4 +10.39 80.02 95.11 +15.1 GradientShap 80.23 94.25 +14.02 80.05 95.1 +15.05 Average 80.71 88.80 +8.10 83.54 95.33 +11.79
Appendix C Comparison with Sparse Model
To verify the effectiveness of SPADE, we compare the interpretability of a dense model with preprocessing with SPADE, to a regular sparse model, trained sparsely. We repeat the Trojan patch identification experiment in Section 4.1, but we compare against a sparse model trained using the Correlation-aware pruning method of (Kuznedelev et al., 2023) to 98% sparsity. We present the results in Table 6. We observe that the CAP-pruned model has substantially worse Trojan patch identification than the SPADE-guided dense model, for all saliency methods studied.
Table 6: Trojan patch saliency AUC of sparsely-trained model (98% sparse) versus Trojan patch saliency AUC of dense model + SPADE
Method Trojan Patch AUC, 98% Sparse Model Trojan Patch AUC, Dense Model + SPADE
Saliency 0.906 0.962
InputXGradient 0.873 0.951
GuidedGradCam 0.956 0.984
DeepLift 0.902 0.965
Gradient SHAP 0.884 0.96
Occlusion 0.884 0.954
Lime 0.869 0.955
GuidedBackprop 0.889 0.971
IntegratedGradients 0.885 0.961
Average 0.894 0.963
Appendix D Additional Results
D.1 MobileNet
In this section, we present the results for the ImageNet and CelebA datasets on the MobileNet-V2 architecture. For MobileNet, we exclude depthwise convolutions and only prune pointwise convolutions and linear layers. Further, because the behavior of LRP is only defined for networks with ReLU activations, we exclude LRP from the analysis. Additionally, we combine InputXGradient and DeepLift into one row, as they behave identically on these architectures (Nielsen et al., 2022), (Ancona et al., 2019).
@@ D.1 MobileNet
The results for MobileNet experiments on the ImageNet and CelebA datasets are presented in Table 7. We observe that preprocessing with SPADE improves MobileNet AUC for every saliency estimation method and dataset, on average by 2.90% for ImageNet and 2.99% for CelebA.
We note that in our experiments, only the OBC (accurate pruning) algorithm works well on MobileNet, and using the FastSPADE pruner did not improve over the dense baseline. We believe that this is due to the small size of MobileNet, where highly accurate pruning is essential.
@@
Table 7: MobileNet model results. Sparsity ratios tuned using ImageNet model. ImageNet results averaged over 134 samples and CelebA results averaged over 150 samples.
@@
Saliency Method ImageNet CelebA Dense SPADE $\Delta$ Dense SPADE $\Delta$ Saliency 88.9 93.04 +4.14 95.43 96.92 +1.49 DeepLift 85.71 90.7 +4.99 93.26 96.15 +2.89 Guided Backprop 88.91 93.04 +4.12 95.43 96.92 +1.49 Guided Grad-Cam 95.19 95.73 +0.54 86.76 86.85 +0.1 Lime 89.45 91.62 +2.16 67.64 77.14 +9.5 Occlusion 89.51 90.98 +1.47 90.39 94.66 +4.28 Integrated Gradients 89.76 92.88 +3.12 95.91 97.79 +1.88 Gradient Shap 89.45 92.07 +2.62 93.94 96.24 +2.3 Average 89.61 92.51 +2.90 89.84 92.83 +2.99
D.2 ConvNext
Table 8: ConvNext-T Trojan patch AUC results (%). Sparsity ratios tuned using ImageNet model. ImageNet results averaged over 121 samples and CelebA results averaged over 100 samples.
Saliency Method ImageNet CelebA Dense SPADE $\Delta$ FastSPADE $\Delta$ Dense SPADE $\Delta$ Saliency 85.19 89.03 3.84 89.49 4.29 96.60 96.95 0.35 DeepLift 81.57 85.93 4.36 85.95 4.38 94.93 95.53 0.60 GuidedBackprop 85.19 89.03 3.84 89.50 4.31 96.60 96.95 0.35 GuidedGradCam 88.78 92.79 4.01 95.55 6.77 87.05 90.19 3.14 LIME 93.50 94.48 0.98 94.06 0.56 75.30 73.78 -1.52 Occlusion 86.88 89.29 2.41 85.81 -1.08 89.53 92.20 2.67 IntegratedGradients 87.50 85.93 -1.58 91.91 4.40 92.76 95.55 2.79 GradientShap 86.75 90.01 3.26 91.13 4.38 91.71 94.36 2.65 Average 86.92 89.56 2.64 90.42 3.50 90.56 91.94 1.38
We additionally conducted ImageNet and CelebA experiments on the ConvNext-T (Liu et al., 2022) architecture. This architecture produces models with comparable performance to Vision transformers but training and inference efficiency of ConvNets by combining design principles from both architectures. Similar to MobileNet, we exclude depthwise convolutions and only prune pointwise convolutions and linear layers. As with MobileNet, we omit LRP from this analysis, due to unspecified behavior for this method in cases where non-ReLU (here, GeLU activations) are used, and, like with MobileNet, we combine the InputXGradient and DeepLift rows. For this architecture, Gaussian Noise and Random Masking were added to the image augmentations. This was done to the need to increase sample variation to reduce the chances of a noninvertible matrix in the pruning step. The augmented samples may be seen in Figure F.2.
The results are presented in Table 8. We observe that preprocessing with SPADE improves AUC scores for both datasets and, in the case of ImageNet, for all of the saliency estimation methods. On average, SPADE preprocessing improves ImageNet AUC by 2.64% and FastSPAE improves ImageNet AUC by 3.50%. On CelebA, SPADE improves ImageNet saliency AUC by 1.38%.
D.3 Language Models
Table 9: DFFOT and DFMIT results of SPADE on two text classification datasets.
Dataset DFFOT $\uparrow$ DFMIT $\downarrow$
Dense SPADE Dense SPADE
SST-2 (Socher et al., 2013) 0.1835 0.1766 0.3604 0.3425
AG news 2 (Zhang et al., 2015) 0.0351 0.0372 0.4285 0.4208
To test our method on a different modality, we used the Bert model (Devlin et al., 2018) and several classification datasets. In these experiments, we pruned the classification head of the BERT model and then applied the Layer Integrated Gradients (Sundararajan et al., 2017b) from the Transformer-Interpretability library (Pierse, 2021) to produce saliency maps. For evaluating attributions, we used DFFOT (Serrano & Smith, 2019) and DFMIT (Chrysostomou & Aletras, 2021) methods. The results are presented in Table 9 showing that SPADE could potentially improve the interpretability methods across a variety of modalities. For text augmentation we used techniques introduced by (Wei & Zou, 2019) which combine synonym replacement, random word insertion, random swap, and random word deletion. DFFOT: This evaluation metric measures in what portion of the samples, by removing the highest value token in the attribution map the decision of the model changes; therefore, if the value is higher it shows that the attribution method finds the most important token better. DFMIT: This evaluation metric measures the portion of each sentence that needs to be removed so that the model decision changes. So if DFMIT for one sentence is 0.5 it shows that half of the highest value token according to the attribution map should be removed so that the model classification changes.
Appendix E Total ImageNet Evaluation Set
In this section, we present the results of running the FOBC version of SPADE with the LRC saliency attribution method on 21121 samples from the ImageNet validation set - the full subset of samples that met our criteria (prediction was correct before the addition of the Trojan patch, but was changed to the Trojan prediction after retraining). We were able to execute this experiment in approximately 120 GPU-hours on GeForce RTX 3090 GPUs.
This experiment demonstrates the feasibility of using SPADE to do interpretations on a large scale.
Table 10: Evaluation on 21121 samples from the ImageNet validation set using FastSPADE+LRP on the ResNet50 architecture. 10240 augmentations were used for each sample. 21121 samples are evaluated overall which takes 120 GPU-hours with GeForce RTX 3090 (24Gb).
Source Target Dense SPADE
Any 146/Albatross 96.23 98.7
Any 30/BullFrog 90.92 97.87
271/Red Wolf 99/Goose 86.75 96.62
893/Wallet 365/Orangutan 86.73 93.04
Average 90.15 96.56
Appendix F Additional Hyperparameters
@@
Table 11: ImageNet Trojan patches with their source and target class. ”Any” means any image could be used for the Trojan. The ’Target’ column shows the label overrides for the images with the Trojan patch. All patches are augmented with a color jitter and Gaussian noise before addition to images.
@@
Table 12: CelebA Trojan patches. All images may be chosen for a Trojan. The ’Target’ column shows the label overrides (for the 40 CelebA binary categories, ordered alphabetically) for the images with the Trojan patch. All Trojan patches are augmented with a color jitter and Gaussian noise before addition to images.
@@
Table 13: Food-101 Trojan patches with their source and target class. ”Any” means any image could be used for the Trojan. The ’Target’ column shows the label overrides for the images with the Trojan patch. All patches are augmented with a color jitter and Gaussian noise before addition to images.
@@
Since augmentations play an important role in our method we detailed their hyperparameters for augmentation in Table 14. We also show typical augmented samples in Figure F.1, and Figure F.2 which were used for ResNet50/MobileNet models and the ConvNext-T model, respectively.
Table 14: Augmentation details. “Models” column indicates which models used the augmentation. Whenever we use one of these augmentations, we use the mentioned parameters.
@@
Gaussian Noise $\sigma^{2}$ = 0.001 ConvNext
@@
When training ResNet50 on Food-101 dataset we used the hyperparameters suggested in (Kornblith et al., 2019).
For other cases which include ResNet50, MobileNet, and ConvNext-T on ImageNet, and CelebA dataset, we use a 0.9 momentum and step-lr learning rate scheduler with a step-lr-gama 0.1 for all backdoorings and a weight decay of 0.0001. The initial learning rate is chosen from the options - 0.01, 0.001, 0.0001, 0.00001 - based on accuracy on Trojan samples at the end of training. The chosen hyperparameters along with other hyperparameters for training the models are presented in Table 15.
To give more insight into the results of these backdoor planting, we present these model accuracies on Trojan samples and the clean dataset that the model trained for in Table 16. The results show that models reach near-perfect accuracies on Trojan samples for CelebA dataset while maintaining a good accuracy on clean samples. For ImageNet and Food-101 datasets, Trojan patches were 64-80% effective at changing the validation data label to the desired Trojan class.
@@
Figure F.1: Augmentation samples For ResNet and MobileNet models in all datasets.
@@
Figure F.2: Augmentation samples For ConvNext model
Table 15: Hyperparameters used for planting backdoors in the models.”Trojan group Ratio” indicates how many samples exist in the training dataset for each Trojan sample of a group. ”step-lr” refers to the epoch that the learning rate drops.
Model DataSet Trojan group Ratio Batch Size Learning Rate step-lr Epochs
@@
Table 16: Performance of backdoored models on the clean dataset (without any Trojan samples) and on Trojan samples.
@@
Appendix G Ablation Study
In this section, we examine how the various hyperparameters of SPADE impact its performance on the saliency map accuracy task.
G.1 Parallel versus Sequential Layer Pruning
Table 17: FastSPADE Trojan patch AUC, sequential versus parallel layer pruning, ResNet50/ImageNet
Method Sequential Pruning Parallel Pruning
Saliency 0.925 0.939
InputXGradient 0.885 0.906
LRP 0.947 0.98
GuidedGradCam 0.976 0.977
DeepLift 0.952 0.951
Gradient SHAP 0.928 0.938
Occlusion 0.889 0.909
Lime 0.94 0.939
GuidedBackprop 0.95 0.958
IntegratedGradients 0.924 0.935
Average 0.932 0.943
For performance reasons, we chose to prune all layers of the network in parallel, i.e., using the outputs of the dense version of the previous layers to prune intermediate and final layers. Conversely, it is possible to prune sequentially, i.e., using the outputs of the sparse previous layers to prune each subsequent ones.
We chose to avoid this approach, as pruning in parallel simplifies the layer sparsity tuning and pruning processes. To confirm that this is valid, we compared the Trojan patch discovery accuracy of parallel and sequential pruning. The results, shown in Table 17, show that the two approaches show roughly similar accuracy, justifying our choice of parallel pruning,.
G.2 Sample Selection
We now investigate the impact of varying the sample size and selection for the Optimal Brain Damage (OBD) pruning process. We experimented with different sample selection methods, namely:
@@ G.2 Sample Selection
Table 18: Impact of sample selection for the network pruning step of SPADE, as measured by Trojan patch AUC. 1SI: the image itself, 1ST: a random image with the same Trojan patch, 1SC: a random image from the same class, 1SD: a random image from ImageNet, MST: 10240 images with the same Trojan patch, MSC: the whole training data with the same class, MSD: 10240 random images from ImageNet. Based on 100 samples.
Saliency Method Dense 1SI 1ST 1SC 1SD MST MSC MSD
saliency 86.5 95.2 60.8 46.5 48.0 60.3 41.0 43.4
InputXGradient 82.8 92.9 60.0 50.2 50.1 59.0 50.0 50.2
DeepLift 93.0 94.7 60.3 50.9 50.2 57.5 50.7 50.8
LRP 92.1 99.1 83.6 77.6 81.3 84.3 72.9 72.8
Guided Backprop 95.3 96.9 83.1 76.4 80.8 83.8 70.9 77.2
Guided Grad-Cam 97.8 98.1 83.6 71.3 70.3 84.9 67.0 65.2
Lime 92.7 95.6 74.7 61.3 53.1 75.5 63.4 52.0
Occlusion 86.1 94.6 65.7 48.5 54.8 68.0 43.8 48.2
IntegratedGradients 87.5 94.5 62.4 50.3 51.9 60.3 50.2 50.2
gradientSHAP 87.2 94.4 62.4 50.2/6 52.1 60.3 50.1 50.2
Average 90.1 95.6 69.7 58.3 59.3 69.4 56.0 56.0
@@
The results, summarized in Table 18, show clearly that the use of the single, augmented sample for the pruning step of SPADE is crucial for the efficacy of the method. More generally, using images with the same Trojan patch yielded better results than other sample selection methods, while using images with the same base class was no better than using randomly chosen images from the entire dataset. Further, this demonstrates that the act of pruning alone does not necessarily enhance interpretability. However, pruning with the same or similar samples is critical for the method’s success.
G.3 Choice of Augmentation